OpenStudy (anonymous):

i need to use reasoning to answer this and also need to find x intercept and y intercept and maxium and minimum please help f(x)=x(x+6)

OpenStudy (anonymous):

this one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

to find x intercept , you set y=0 and solve for x y=x(x+6) 0=x(x+6) x=0,-6 to find y intercept , set x=0 and solve for y

OpenStudy (anonymous):

to find min/max, find vertex using following formula -b/2a

OpenStudy (anonymous):

so i confused on how to solve for y

OpenStudy (anonymous):

i dont know how to do any of this my teacher just gave it to me and told me to do it